OpenStudy (anonymous):

what would be a good situation to make out of 5x+y=0 and 5x+2y=30

OpenStudy (anonymous):

5x+y=0 -> y=-5x From the second equation: 5x+2(-5x)=30 Solve the above for x. Can you carry of from here?

OpenStudy (anonymous):

i need more of a word problem

OpenStudy (anonymous):

An english math statement which can be symbolized into the expressions, 5x+y=0 and 5x+2y=30 ?

OpenStudy (anonymous):

yes please i just cant think of anything

OpenStudy (anonymous):

Frankly, neither can I.

OpenStudy (anonymous):

man its okay thanks for the help